OpenStudy (anonymous):

a man is currently 20 years older than hs oldest son. his oldest son is currently 2 years older than his youngest son. 3 years from now, the product of the ages of his sons will be ten more than eleven tiems the age of the father. what is the current age of the mans youngest son

OpenStudy (asnaseer):

lets call the fathers age 'f', the eldest sons age 'e' and the youngest sons age 'y'. then you are give: f = 20 + e e = 2 + y (e+3)(y+3) = 10 + 11(f+3) so you now have 3 equations with 3 unknowns that you should be able to solve.